2. Two resistors R and R2 may be connected either in series or in parallel across a battery with emfe. The power consumed by the parallel combination is five times that of the series combination. Find the ratio R R2/ (RR2)
Answer #1

power consumed=voltage^2/total resistance

so power consumed in parallel configuration / power consumed in series configuration=total resistance in series configuration / total resistance in parallel configuration

==>5=(R1+R2)/(R1*R2/(R1+R2))

==>(R1+R2)^2=5*R1*R2

==>R1R2/(R1+R2)^2=1/5